OpenStudy (anonymous):

find the vertex, line of symmetry, the maximum or minumum value of the quadratic funciton f(x)=-2x^2+2x+3

OpenStudy (anonymous):

f'(x) = -4x + 2 = 2(-2x+1) x= 1/2 y = 3.5 (7/2)

OpenStudy (anonymous):

vertex = (1/2,7/2) LoS = x = 1/2 max = y = 7/2

OpenStudy (anonymous):

I have already posted the solution in your first thread devette

OpenStudy (anonymous):

Sub x=1/2 into the eqn again raheen